This question is typical on some driver’s license exams: A car moving at 43 km/h skids
16 m with locked brakes.
How far will the car skid with locked brakes
at 86 km/h? Assume that energy loss is due
only to sliding friction.
Answer in units of m
1
Expert's answer
2020-12-23T06:54:29-0500

"s_1=\frac{v_1^2-v_0^2}{2a}\to a=-\frac{v_0^2}{2s}=-\frac{11.94^2}{2\cdot16}=-4.46(m/s^2)"


"s_2=\frac{v_2^2-v_0^2}{2a}=\frac{0^2-23.89^2}{2\cdot(-4.46)}\approx64(m)"